# Breaking the News: A LLM-based Game where Players Act as Influencer or Debunker for Raising Awareness About Misinformation

### v1

@@ 5.2.1. Raise Awareness of Misinformation though gameplay
In-game News as a Reflection of Real-world Misinformation 23 out of 47 participants noted that the in-game news mirrored real-world situations, thereby heightening their awareness of the characteristics of misinformation. A common observation was that news is rarely entirely true or false; instead, it often presents a mixture of both. This complexity makes genuine misinformation more challenging to detect. As one participant stated:
@@
Participants also learned from observing their opponents. For example, N22, who played the role of a debunker, noticed how the misinformation creator crafted and disseminated false information to persuade others:
@@
5.2.2. Identifying Misinformation through Source Evaluation:
After the gameplay sessions, participants reported increased awareness of the varying credibility of different information sources. The game helped them to realize that producers of misinformation often seeks to enhance credibility by deliberately referencing authoritative organizations. One participant reflected on this realization:
5.2.3. Identifying Misinformation through Emotional Manipulation Tactics
Participants learned various tactics for both creating and debunking misinformation through the game’s instructions and their in-game experiences. A particularly commonly identified tactic was emotional manipulation, which was noted by 35 out of 47 participants (18 misinformation creators and 17 debunkers). By analyzing the game logs, we identified common emotional manipulation strategies used in the game. Most players crafted messages designed to evoke anxiety and fear, while some also attempted to generate feelings of hope. For example, in Round 4, Player 1 spread rumors about a doctor’s death, which directly incited public panic (N1). Across all rounds, Player 1 frequently used emotional appeals and personal stories to enhance the perceived credibility of the misinformation (N6, N8). Additionally, invoking cultural pride and heritage was a powerful tactic used to build trust in misinformation (N4), while celebrity endorsements further increased the complexity and believability of the misinformation (N9). As illustrated in Figure 8, players significantly increased public trust in their information by using emotionally charged language (N31). These strategies align closely with the characteristics of misinformation, where emotional appeals are commonly used to influence public opinion (Chuai et al., 2022).
Interestingly, when playing the role of the debunker, participant can reflected on the emotionally inflammatory language used by the misinformation creator and helped them develop a more clear strategy for addressing misinformation. This approach involved separating the factual content of a message from its emotional manipulations and focusing more on the factual aspects, as one participants explained
5.2.4. Critical Thinking about Misinformation Motives
After the gameplay sessions, many participants (31 out of 47) demonstrated an enhanced understanding of the intricacy of information and the varied perspectives it can convey. This experience increased their awareness of the importance of considering the motivations behind messages. Participants also mentioned that they are now more inclined to think critically about the goals behind the information they encounter, especially in real-life situations where such considerations are common. One participant explained how the game illustrated the pre-determined nature of many messages:
5.2.5. The Impact of Role-Playing as a Misinformation Creator on Learning
Through the experience of playing the role of the misinformation creator (Player 1), some participants became aware of just how low the barriers are for creating misinformation. This made them more cautious about the influence of certain public figures, particularly online ”influencers.” As one participant noted:
N45: I’m Player 1, and I realized that the cost of creating rumors is so low. If I were an online celebrity or someone with the ability to influence public opinion, and my job wasn’t that of a journalist, I might not need to be very responsible for spreading these kinds of rumors.
Another participant reflected on how playing as a misinformation creator broadened their perception of misinformation, particularly regarding how easily false information can be fabricated. This experience expanded their understanding of the boundaries of misinformation, making them more aware of how easily those boundaries can be crossed:
@@ 5.2.5. The Impact of Role-Playing as a Misinformation Creator on Learning
These findings indicate that when participants took on the role of distributing misinformation, it helped them to better grasp how misinformation is produced and emphasized how easily it can spread. This is further proved in the game. As shown in Figure 9, Player 1 employed certain strategies when faced with an unfavorable context, such as avoiding or distorting facts and creating a positive image. In the interview, Player N31 also reported that as the game progressed, they felt increasingly confident in their ability to generate misinformation.
@@
5.2.6. Tailoring Debunking Strategies to Audience Characteristics
Many participants found the responses of the LLM-simulated personas to be particularly engaging. They analyzed these responses to understand the reasons behind changes in opinion, how the output of other players influenced these shifts, and what the personas now trusted. Participants noted that the LLM-simulated personas provided clear trust level scores and reactions, which were helpful in organizing their responses. As one participant observed:
@@ 5.2.6. Tailoring Debunking Strategies to Audience Characteristics
This insight into the personas’ dynamic responses helped players refine their strategies. For example, players noticed that different personas reacted differently to emotional and logical appeals. While three personas were easily swayed by emotional arguments, the other two preferred rational, science-based evidence. Recognizing these tendencies, Player 2 successfully countered emotional tactics through logical analysis and evidence (N35). However, even when Player 2 consistently employed logical reasoning and evidence throughout the rounds, it didn’t always succeed in shifting all five personas’ trust levels in their favor. Participants further realized that using evidence to dispute misinformation wasn’t always effective. For instance, one persona was a traditional-minded person who resisted new scientific findings. As shown in Figure 10, after several rounds of gameplay, players adapted their strategies to persuade the persona by considering her perspective. In the follow-up interview, a participant reflected:
@@
5.2.7. Improving Confidence in Debunking
Many participants reported an increase in their confidence to debunk misinformation in real-life scenarios. The positive outcomes they experienced in the game appeared to strengthen their ability to challenge misinformation outside the game environment. For example, in Round 3, N7 purchased a hint to access supporting data for their argument. However, N8 used the same data to identify flaws and construct a counterargument. By leveraging data-driven reasoning, N8 effectively challenged N7’s claims and strengthened their own position, as shown in Figure 11. In the post-game interview, N8 reported their increased confidence in using evidence-based rebuttals:
N8: I’d say my confidence went up to around 70% or 80%, because I realized that presenting evidence and discuss this is really effective.
However, some participants expressed concern that debunking misinformation in real life would be more challenging. Unlike in the game, where the personas’ cognitive levels and trust levels were explicitly shown, it is much harder to gauge these factors in real-world interactions. This uncertainty could make it more difficult to challenge misinformation effectively in everyday situations.
Figure 11. An example of how a debunker identifies logical flaws in the information by comparing different pieces of information.
5.2.8. Impact of Gameplay on Future Debunking Actions
A few participants (4/47) shared that the game increased their likelihood of taking action against misinformation in the future. This change in attitude was driven either by participants’ previous negative personal experiences with misinformation or by their realization during gameplay of the serious consequences misinformation can have. The gameplay experience enhanced their willingness to invest time and effort into distinguishing and debunking false information. As one participant said:
5.2.9. Challenges and Negative Effects of Gameplay
While some participants gained confidence in their ability to debunk misinformation, others experienced a decrease in confidence (5/47). These participants observed that the game was close to real life, particularly that some individuals held strong pre-existing beliefs that were difficult to challenge. The game reinforced this reality, leading to a reduction in their confidence. As one participant noted:
Another challenge reported was the overwhelming amount of information presented in the game. In each round, players had to process news stories, the reactions of five personas and their changes, as well as their opponent’s output. After absorbing this information, participants were required to devise strategies and craft tailored responses. The volume of information, particularly by the end of the game, left some participants feeling exhausted, which may have impacted their performance. In addition, the game’s mechanics required players to have a basic level of media knowledge to effectively take on the roles of “influencer” and “debunker.” Some participants also noted that the competitive nature of the game could lead to an unbalanced experience if one player was significantly stronger than the other. As one participant noted:
N40: The game was very fun, then I felt a bit nervous, and by the time I got to the end, it was a bit exhausting. I felt nervous because there was a lot of information at the beginning, and I was competitive with Player 2. There was much writing involved, and I felt uncertain because I’m not very good at writing, and I knew Player 2 was very skilled. So I felt a little nervous.
@@ 6.1. General Discussion
Our study provided empirical evidence that the intervention improved some aspects of users’ general media literacy, their behavioral intentions to use different strategies for checking misinformation and their performance in discriminating between fake and real news. However, it did not substantially enhance players’ prosuming skills, despite the mechanics of the game being tailored to support content generation. One possible explanation of this result would be that pre-survey data indicated that participants entered the study with high levels of confidence in their prosuming abilities. This ”ceiling effect” likely constrained measurable gains, as participants had or perceived less room for improvement. Another explanation is that high-level skills of critical prosuming require collaborative efforts and collective intelligence between participants (Lin et al., 2013), which was not possible to achieve in the PvP model of the game. Finally, the game only provided four rounds of content creation for each person, which could be sufficient to apply a critical perspective on the content, but not enough to train creation effectiveness.
@@
Another important finding is that, similarly to the previous studies of another game-based intervention (Leder et al., 2024) we did not find significant effects from the intervention on the person’s self-confidence in tackling misinformation. Interestingly, game log data indicated that players’ proficiency improved during gameplay; by the third and fourth rounds, they typically produced longer and more comprehensive messages compared to the initial rounds. Social Cognitive Theory (Bandura, 1997). may explain this discrepancy. According to it, enactive mastery experiences — ie., successful task completion— are the most influential sources of self-efficacy(Bandura, 1997). Positive experiences bolster self-efficacy, while repeated failures undermine it. In our game, success was determined by evaluations from LLM-simulated public opinion. The qualitative results revealed that the players found that no single strategy was effective for all characters, requiring frequent adjustments based on the unique characteristics of each character. This unpredictability made it difficult to achieve consistent success. While players developed greater proficiency during the game, the difficulty in achieving consistent success may have limited their perceived self-efficacy. However, this outcom can also be considered through the lens of the educational effects of the game. Previous studies about differential challenges of misinformation showed that young adults often overestimate their ability to assess information effectively(Papapicco et al., 2022; Porat et al., 2018); in this context, the reaction of our participants, most of whom were young adults could be a positive signal that they became aware of the complexity of misinformation and the absence of one-size-fit-all solutions.
@@
Lastly, In our qualitative findings, participants reported skepticism toward information that purely relied on authoritative sources because they found their opponents used fabricated evidence from these sources to gain the trust of the LLM-simulated characters, with a view to increasing their game score. This gameplay experience reminded them of real-life situations, where misinformation often exploits trust by citing credible authorities. As a result, players learned to examine the intent behind messages instead of automatically trusting authoritative sources. This evolution in how participants evaluated information aligns with the Elaboration Likelihood Model (Petty et al., 1984), which describes how individuals process persuasive messages via central (critical evaluation of content) or peripheral routes (reliance on heuristic cues). In the gameplay, participants appeared to shift from peripheral processing (trusting authority as a shortcut) to central processing (evaluating the intent and content of the message) when exposed to the misuse of authority in gameplay. Furthermore, our findings align with studies showing that gamified inoculation techniques for pre-bunking misinformation can trigger skepticism related to both false and real news(Hameleers, 2023; Modirrousta-Galian and Higham, 2023). While this induced skepticism might seem limiting —potentially thus undermining trust in high-credibility sources— we believe it supports the goal of fostering critical media literacy. By encouraging players to evaluate the trustfulness of content, source and intent, the game develops essential skills for navigating today’s complex information landscape. Rather than promoting cynicism, this skepticism cultivates constructive inquiry that helps individuals better discern reliable information. Notably, many participants pointed to the need to cross-check sources as a practical solution for verifying information, demonstrating their enhanced critical thinking skills.
6.2. Outcomes of the Game Mechanics
6.2.1. PvP model for Media-Literacy Game
@@
Unlike prior misinformation education game which broadly employ single-player mechanics (Roozenbeek and Van der Linden, 2019; Lewsey, 2020; har, 2020; Jeon et al., 2021); ”Breaking the News” applies PvP mechanics. While in general, previous studies in serious games showed that PvP games are more engaging and motivating (Cagiltay et al., 2015); at the same time in competitive environments, the motivation to “win” may overshadow educational goals. In addition, in recent studies of another misinformation game it showed that some students may not find competition enjoyable or motivating(Axelsson et al., 2024). Our results demonstrated that, in general, our participants were highly motivated to play one against another. Moreover, they indicated that they learned from each other’s strategies, and it helped them better understand the dynamic of misinformation. Therefore, we can conclude that, in our case, this approach was beneficial to fulfilling the intended purpose of the game. However, our study was conducted on East Asian participants, who are part of a collectivist-oriented rather than competitive culture (Chung and Mallery, 1999). Previous studies have shown that cultural factors play an important role in the degree of competitiveness in gamified interventions (Oyibo et al., 2017). Therefore, it is also possible that in other cultural settings, the game’s incentives can trigger more intense competition, which can negatively affect educational results.
@@ 6.2.2. Free-form Input Generation
One notable feature of Breaking the News is its free-form response format, which contrasts with the linear choice-based formats commonly used in prior game intervention(Roozenbeek and Van der Linden, 2019; Lewsey, 2020; har, 2020; Jeon et al., 2021; Micallef et al., 2021). Previous works in the field of role-playing games showed that this approach makes game interaction more natural and increases user engagement (Csepregi, 2021; Ashby et al., 2023). Based on our data, we can say that this format positively influences players’ learning outcomes by enhancing engagement and replayability. By allowing players to develop narratives themselves, each playthrough feels unique, encouraging players to return to the game. This variability not only boosts intrinsic motivation but also contributes to the game’s educational value. When players repeatedly engage with core mechanics through varied narratives, they continue practicing essential skills, thereby reinforcing and enhancing their learning outcomes(Kücklich and Fellow, 2004).
In an educational setting, previous works showed that in choice-based formats, interactions are typically brief, requiring students to select predetermined options. This structure could lead to ”guesswork,” where learners may choose correct answers without fully understanding the underlying concepts. In contrast, free-form responses force students to actively reason and articulate original ideas, fostering deeper engagement and critical thinking (Bryfczynski, 2012). This autonomy allows players to craft responses based on their understanding of their roles, making this another motivating factor for returning to the game(Ravyse et al., 2017).
@@ 6.2.3. LLM-Powered Feedback
Another innovation of Breaking the News lies in incorporating LLM-based ”characters” as evaluators, creating interactive feedback for players. Although there are other attempts to incorporate LLMs into helping users learn about misinformation (Danry et al., 2023; Hsu et al., 2024) (including gamified attempts(Tang and Singha, 2024)), in these approaches, AI was mostly the source of correcting information. While this approach has benefits, it is can also be criticised due to the possibility of LLMs creating incorrect by plausible text (Kim et al., 2024; Agarwal et al., 2024). In this case, it is possible that the intervention will disinform people to an even greater extent. In contrast, in our study, we incorporate AI not as a source of information but as non-playable characters with their own opinions. Therefore, the educational part of the game becomes more robust to resist the negative effects of erroneous generations, as they can only affect the opinion of the ”character” but not the main narrative of the game. In general, LLMs demonstrate the ability to simulate human behavior and reactions, consistent with findings from prior research(Park et al., 2023b); we also find, that the system is capable of emulating the opinion of 5 different characters at the same time and still keeping them consistently different during the game.
One advantage of dynamic LLM feedback is that it achieves greater engagement compared to binary feedback (e.g., true or false). Based on our observation, players adapted by employing alternative persuasive strategies tailored to the character and concentrated on the character’s feedback. Notably, players exhibit higher levels of engagement with characters they found relatable due to personal experiences, focusing more attentively on their feedback. This level of engagement is difficult to achieve with traditional binary feedback mechanisms, which provide limited insights beyond checking their correctness.
6.3. Design Implications for the Further Development of Serious Games with LLM-components
Based on our experience designing and testing the game, we summarize the following recommendations to practitioners working on similar projects.

### v2

@@ 5.2. RQ2: Player Behaviors in Response to Game Mechanics and Opponent Tactics
In-game News as a Reflection of Real-world Misinformation: 23 out of 47 participants noted that the in-game news mirrored real-world situations, thereby heightening their awareness of the characteristics of misinformation. A common observation was that news is rarely entirely true or false; instead, it often presents a mixture of both. This complexity makes genuine misinformation more challenging to detect. As one participant stated:
@@
Participants also learned from observing their opponents. For example, N22, who played the role of a Journalist, noticed how the Influencers crafted and disseminated false information to persuade others:
@@
The Impact of Role-Playing as a Influencer on Learning: Through the experience of playing the role of the Influencer, some participants became aware of just how low the barriers are for creating misinformation. This made them more cautious about the influence of certain public figures, particularly online Influencers. As one participant noted:
N45: I’m Influencer, and I realized that the cost of creating rumors is so low. If I were an online celebrity or someone with the ability to influence public opinion, and my job wasn’t that of a Journalist, I might not need to be very responsible for spreading these kinds of rumors.
Another participant reflected on how playing as an Influencer broadened their perception of misinformation, particularly regarding how easily false information can be fabricated. This experience expanded their understanding of the boundaries of misinformation, making them more aware of how easily those boundaries can be crossed:
@@
These findings indicate that when participants took on the role of distributing misinformation, it helped them to better grasp how misinformation is produced and emphasized how easily it can spread. This is further proved in the game. As shown in Figure 8, Influencer employed certain strategies when faced with an unfavorable context, such as avoiding or distorting facts and creating a positive image. In the interview, Participant N31 also reported that as the game progressed, they felt increasingly confident in their ability to generate misinformation.
@@
Tailoring Debunking Strategies to Audience Characteristics: Many participants found the responses of the LLM-simulated characters to be particularly engaging. They analyzed these responses to understand the reasons behind changes in opinion, how the output of other players influenced these shifts, and what the characters now trusted. Participants noted that the LLM-simulated characters provided clear trust level scores and reactions, which were helpful in organizing their responses. As one participant observed:
@@
This insight into the characters’ dynamic responses helped players refine their strategies. For example, players noticed that different characters reacted differently to emotional and logical appeals. While three characters were easily swayed by emotional arguments, the other two preferred rational, science-based evidence. Recognizing these tendencies, The Journalist successfully countered emotional tactics through logical analysis and evidence (N35). However, even when consistently employed logical reasoning and evidence throughout the rounds, it didn’t always succeed in shifting all five characters’ trust levels in their favor. Participants further realized that using evidence to dispute misinformation wasn’t always effective. For instance, one persona was a traditional-minded person who resisted new scientific findings. As shown in Figure 9, after several rounds of gameplay, players adapted their strategies to persuade the persona by considering her perspective. In the follow-up interview, a participant reflected:
@@
Challenges and Negative Effects of Gameplay: While some participants gained confidence in their ability to debunk misinformation, others experienced a decrease in confidence (5/47). These participants observed that the game was close to real life, particularly that some individuals held strong pre-existing beliefs that were difficult to challenge. The game reinforced this reality, leading to a reduction in their confidence. As one participant noted:
Another challenge reported was the overwhelming amount of information presented in the game. In each round, players had to process news stories, the reactions of five characters and their changes, as well as their opponent’s output. After absorbing this information, participants were required to devise strategies and craft tailored responses. The volume of information, particularly by the end of the game, left some participants feeling exhausted, which may have impacted their performance. In addition, the game’s mechanics required players to have a basic level of media knowledge to effectively take on the roles of Influencer and Journalist. Some participants also noted that the competitive nature of the game could lead to an unbalanced experience if one player was significantly stronger than the other. As one participant noted:
N40: The game was very fun, then I felt a bit nervous, and by the time I got to the end, it was a bit exhausting. I felt nervous because there was a lot of information at the beginning, and I was competitive with Journalist. There was much writing involved, and I felt uncertain because I’m not very good at writing, and I knew (my opponent) was very skilled. So I felt a little nervous.
@@ 6.1. Summary and Interpretation Results
Our study provided evidence that the game intervention improved some aspects of users’ media literacy, their intentions to check misinformation, and their ability to distinguish fake from real news. However, it did not significantly enhance prosuming skills, even though the game mechanics were designed to support content creation. One reason for this could be that participants started the study with high confidence in their prosuming abilities, creating a “ceiling effect” that limited measurable improvement. A possible explanation is that high-level prosuming skills require collaborative efforts (Lin et al., 2013), which were not possible in the PvP model of the game. Also, the game only had four rounds of content creation, which might have been enough to apply a critical perspective, but not enough to significantly improve creation effectiveness. Interestingly, both the game intervention and the control intervention resulted in an improvement in the willingness to apply verification behaviours. That means that exposure to labelled misinformation can be a mechanism to orient people towards fact-checking behaviours; however, as we saw in the results of the MIST test, it is not enough to improve the ability to identify it.
@@
Another key finding is that, similar to other game intervention (Leder et al., 2024), we did not find significant effects on participants’ self-confidence in dealing with misinformation. Game log data, however, showed that players’ proficiency improved as they progressed. By the third and fourth rounds, players typically produced longer, more comprehensive messages. This discrepancy can be explained by Social Cognitive Theory (Bandura, 1997). According to the theory, successful task completion are the most powerful sources of self-efficacy (Bandura, 1997). In our game, success was determined by LLM-simulated public opinion. The results revealed that no single strategy worked for all characters. While players developed greater proficiency during the game, the difficulty in achieving consistent success may have limited their perceived self-efficacy. However, this outcome can also be considered through the lens of the educational effects. Previous studies showed that young adults often overestimate their ability to assess information effectively(Papapicco et al., 2022; Porat et al., 2018); in this context, the reaction of our participants, most of whom were young adults could be a positive signal that they became aware of the complexity of misinformation and the absence of one-size-fit-all solutions. The results of the control group support this interpretation. In a control group, we observed that the exposure to misinformation materials positively affected the participants’ confidence in their ability to tackle misinformation; however, it did not improve their skills in discriminating between true and false news. That potentially means that the lower level of confidence in tackling the issue can be a good outcome, showing a more cautious and reflective approach to the issue.
@@
Lastly, participants reported being skeptical of information relying solely on authoritative sources. They observed that opponents used fabricated evidence from these sources to gain trust and improve their game scores. This gameplay experience reminded them of real-life situations, where misinformation often exploits trust by citing credible authorities. As a result, players learned to examine the intent behind messages instead of automatically trusting authoritative sources. This shift aligns with the Elaboration Likelihood Model (Petty et al., 1984), which describes how individuals process persuasive messages via central (critical evaluation of content) or peripheral routes (reliance on heuristic cues). In the gameplay, participants appeared to shift from peripheral processing (trusting authority as a shortcut) to central processing (evaluating the intent and content of the message) when exposed to the misuse of authority in gameplay. Our findings also align with studies showing that gamified inoculation techniques can increase skepticism toward both false and real news (Hameleers, 2023; Modirrousta-Galian and Higham, 2023). While this skepticism may seem to undermine trust in high-credibility sources, it supports the goal of developing critical media literacy. By encouraging players to evaluate content, sources, and intent, the game helps them navigate today’s complex information landscape. Rather than fostering cynicism, this skepticism encourages inquiry, aiding in the discernment of reliable information. Notably, many participants also suggested cross-checking sources as a practical solution, demonstrating their enhanced critical thinking skills.
6.2. Game Mechanics and Learning Outcomes
6.2.1. PvP model for Media Literacy Game
Unlike prior game interventions that generally use single-player mechanics (Roozenbeek and Van der Linden, 2019; Lewsey, 2020; Roozenbeek and van der Linden, 2020; Jeon et al., 2021), Breaking the News uses PvP mechanics. While previous studies have shown that PvP games are often more engaging and motivating (Cagiltay et al., 2015), the competitive environments where the motivation to “win” may overshadow educational goals. In addition, recent research found that some students did not find competition enjoyable or motivating (Axelsson et al., 2024). In this respect, it is important to be sure that the tested interaction lies within the borders of “constructive competition” - competition which internally feels like working on mutual improvement and therefore raising intrinsic motivation (Sailer and Homner, 2020).
@@ 6.2.1. PvP model for Media Literacy Game
Our results showed that participants were highly motivated to compete against each other. They also indicated that learning from others’ strategies helped them understand misinformation dynamics better. Therefore, in our case, the PvP approach served the intended educational purpose well. However, our study was conducted with East Asian participants, who come from a collectivist culture rather than a competitive one (Chung and Mallery, 1999). Previous studies have suggested that cultural factors significantly affect competitiveness in gamified interventions (Oyibo et al., 2017), so it is also possible that in other cultural settings, the game’s incentives can trigger more intense competition, which can negatively affect educational results.
@@ 6.2.2. Free-form Input Generation
One notable feature of Breaking the News is its free-form response format, which contrasts with the linear choice-based formats commonly used in prior game interventions (Roozenbeek and Van der Linden, 2019; Lewsey, 2020; Roozenbeek and van der Linden, 2020; Jeon et al., 2021; Micallef et al., 2021). Our findings support prior research (Csepregi, 2021; Ashby et al., 2023), suggesting that free-form response formats positively influences players’ learning outcomes by enhancing engagement and replayability. By allowing players to develop narratives themselves, each playthrough feels unique, encouraging players to return to the game. In educational settings, previous studies have shown that choice-based formats typically involve brief interactions where learners select predetermined options. This can lead to “guesswork,” with students selecting correct answers without fully understanding the concepts. In contrast, free-form responses force players to reason and articulate their ideas, fostering deeper engagement and critical thinking (Bryfczynski, 2012). This autonomy allows players to craft responses based on their understanding, providing an additional motivaton to return to the game (Ravyse et al., 2017).
@@ 6.2.3. LLM-Powered Feedback
Another innovation of Breaking the News lies in its interaction and feedback mechanisms. Although other attempts have used LLMs to help users learn about misinformation (danry2023dont; Hsu et al., 2024) (including gamified attempts (Tang and Singha, 2024)), in these approaches, AI was mostly the source of correcting information. While this approach can be useful, it has been criticized for the risk of LLMs generating incorrect but plausible text (Kim et al., 2024; Agarwal et al., 2024). In this case, it is possible that the intervention will disinform people to an even greater extent. In contrast, our study uses AI not as an information source, but as non-playable characters with their own opinions. This design makes the educational aspect of the game more robust against potential errors, as incorrect AI-generated feedback only affects the character’s opinion, not the main narrative. In general, LLMs demonstrate the ability to simulate human behavior and reactions, consistent with fndings from prior research(Park et al., 2023). We also found that the system could emulate the opinions of five distinct characters while maintaining consistency throughout the game.
One advantage of dynamic LLM feedback is that it achieves greater engagement compared to binary feedback (e.g., true or false). Based on our observation, players adapted by employing alternative persuasive strategies tailored to the character and focused on the feedback they received. Notably, players showed higher engagement with characters they could relate to from personal experience, paying more attention to their feedback. This level of engagement is difficult to achieve with traditional binary feedback, which offers limited insights beyond checking their correctness.
6.3. Design Implications for Future Media Literacy Games